10lbs of mixed nuts containing 30% peanuts were mixed with 18lbs of peanuts. Peanuts are what percent of the new mixture?
i tried this:
.3=10+18

Answer by stanbon(75887)

10lbs of mixed nuts containing 30% peanuts were mixed with 18lbs of peanuts. Peanuts are what percent of the new mixture?
======================
Equations:
Let x be the percent of peanuts in the mixture.
peanuts + peanuts = peanuts
0.30*10 + 18 = x(10+18)
3 + 18 = x(28)
x = 21/28
x = 0.75
----
x = 75%
====================
Cheers,
Stan H.